### v1

# WorldSense: Evaluating Real-world Omnimodal Understanding for Multimodal LLMs

Abstract
In this paper, we introduce WorldSense, the first benchmark to assess the multi-modal video understanding, that simultaneously encompasses visual, audio, and text inputs. In contrast to existing benchmarks, our WorldSense has several features: (i) collaboration of omni-modality, we design the evaluation tasks to feature a strong coupling of audio and video, requiring models to effectively utilize the synergistic perception of omni-modality; (ii) diversity of videos and tasks, WorldSense encompasses a diverse collection of 1,662 audio-visual synchronised videos, systematically categorized into 8 primary domains and 67 fine-grained subcategories to cover the broad scenarios, and 3,172 multi-choice QA pairs across 26 distinct tasks to enable the comprehensive evaluation; (iii) high-quality annotations, all the QA pairs are manually labeled by 80 expert annotators with multiple rounds of correction to ensure quality. Based on our WorldSense, we extensively evaluate various state-of-the-art models. The experimental results indicate that existing models face significant challenges in understanding real-world scenarios ( best accuracy). We hope our WorldSense can provide a platform for evaluating the ability in constructing and understanding coherent contexts from omni-modality.
1 Introduction
The ability to comprehend and reason about multimodal inputs—ranging from visual and textual to auditory, tactile, and beyond—is fundamental for both human and artificial agents to navigate and interpret the world. For example, when driving a car, a human driver integrates visual information (e.g., recognizing road signs, traffic lights, and obstacles), auditory cues (e.g., hearing the honking of another car or a siren approaching from behind), and tactile feedback (e.g., the feel of the steering wheel, the vibrations of the road, or the responsiveness of the brakes) to make real-time decisions and ensure safe navigation. This seamless multimodal integration enables intelligent agents to process complex, dynamic environments and respond to subtle cues—an ability that is essential for both human perception and the development of embodied agents designed to interact naturally in the world.
In the recent literature, the development of Multi-modal Large Language Models (MLLMs) [49, 24, 48, 57, 59, 75, 43, 12] have led to remarkable progress on a series of tasks, for example, classification [36], captioning [1, 10, 35], question-answering [56, 50, 40], OCR [45, 76], segmentation [25, 67, 22], autonomous driving [46, 52, 4] and more. However, multi-modal analysis primarily focuses on visual-language information, leaving out crucial modalities like audio, which results in an incomplete evaluation of their multimodal capabilities. While some benchmarks have begun incorporating both visual and audio modalities, they still exhibit several limitations. For example, OmniBench [31] and AV-Odyssey Bench [19] mainly emphasize image evaluation, whereas other benchmarks [18, 29, 70] either restrict to captioning tasks or are limited to simple scenarios, or suffer from low-quality, monotonous questioning patterns.
In this paper, we introduce WorldSense, the first benchmark designed to comprehensively evaluate MLLMs’ ability to perceive, understand, and reason through the integration of omni-modal information in real-world contexts. WorldSense is defined by three key features: (i) Collaboration of Omni-Modality: At the core of WorldSense is the emphasis on the interplay between audio and visual information, as illustrated in Figure 1. Each question is designed to require both modalities for a correct answer; the absence of either modality leads to an incorrect response. Such design rigorously tests a model’s ability to simultaneously process multiple sensory inputs for accurate understanding; (ii) Diversity of Videos and Tasks: WorldSense encompasses 1,662 audio-visual synchronized videos across 8 domains and 67 fine-grained subcategories. It includes 3,172 multiple-choice questions spanning 26 distinct cognitive tasks, from basic perception to complex reasoning. This diversity enables a systematic and comprehensive evaluation of MLLMs’ multimodal understanding capabilities; (iii) High-Quality Annotations: To ensure reliability, all QA pairs in WorldSense are manually labeled by 80 expert annotators and undergo multiple rounds of validation through both human review and automatic MLLM verification. This rigorous process ensures the benchmark’s accuracy and reliability. By integrating these features, WorldSense sets a new standard for evaluating MLLMs’ ability to handle real-world complexity, bridging the gap between artificial intelligence and human-like multimodal understanding.
We conduct comprehensive experiments on WorldSense to evaluate a range of existing MLLMs, including open-source video models, video-audio models, and proprietary models. Our results highlight significant limitations in the current models’ ability to understand and reason through omni-modal information in real-world contexts. First, while open-source video-audio models can process both video and audio inputs, they achieve only around 25% accuracy on our benchmark, comparable to random guessing. In contrast, proprietary models, such as Gemini 1.5 Pro, which processes both video and audio, achieve the highest accuracy at 48%. However, when provided with only a single modality (either video or audio), Gemini 1.5 Pro’s performance drops significantly by approximately 15%, underscoring the crucial role of omni-modal collaboration in real-world tasks. These findings emphasize the strong coupling of modalities in WorldSense and reveal a substantial gap in current MLLMs’ real-world understanding capabilities.
Additionally, we further conduct detailed ablation studies on how different modalities impact model performance. Our analysis shows that visual information plays significant for understanding, while adding audio—whether as raw signals or transcribed subtitles—also improves results. However, raw audio inputs yield superior gains, emphasizing the importance of paralinguistic features such as prosody, intonation, and acoustic context, which are often lost in text transcription. Additionally, increasing the density of video frame sampling, which provides richer temporal information, consistently improves performance. These findings underscore the complementary nature of visual and auditory modalities and highlight the necessity of integrating both for robust real-world understanding.
To summarise, in this paper, we have made three contributions: (i) we introduce, WorldSense, the first benchmark specifically designed to evaluate MLLMs’ omni-modal understanding capabilities in real-world scenarios. WorldSense is characterized by the integration of omni-modality, diverse video categories, and high-quality QA pairs; (ii) We have conducted extensive experiments to evaluate current MLLMs’ capabilities in real-world omni-modal understanding. Our results demonstrate that open-source video-audio MLLMs perform only marginally better than random guessing, while even the best-performing proprietary models achieve only accuracy, revealing significant room for accurate omni-modal understanding; (iii) Through detailed analysis, we investigate the crucial factors affecting real-world omni-modal understanding. Our findings demonstrate that acoustic information, visual cues, and dense temporal sampling all substantially impact model performance. These finds provide valuable guidance for the future development of real-world omni-modal understanding.
2 Related Work
Multimodal Large Language Models. Current Large Language Models (LLMs) are capable of processing multimodal information, including visual, text, and audio. Early works, such as [75, 35, 80, 11, 64, 51], successfully combine vision and text modalities. Subsequent research extends to temporal understanding [66, 24, 58, 39, 61, 65, 26, 14, 69, 74, 60, 6, 41, 34], while parallel efforts [55, 9, 8] focus on audio processing. Recently, researchers shift attention to models [7, 54, 58, 42, 59] capable of simultaneously processing text, vision, and audio inputs. Despite the growing interest in the models which can perform the omnimodality understanding, the absence of a comprehensive evaluation benchmark restricts the development. To address this limitation, we introduce our WorldSense to evaluate models’ capabilities in perceiving and understanding real world omnimodal scenarios.
Multimodal Benchmarks. The development of MLLMs has been driven by benchmarks, evolving from static image understanding [77, 37, 28, 27, 16, 73] to temporal comprehension [30, 38, 53, 79, 13, 17, 23, 63, 68, 72, 33, 3]. However, these benchmarks largely overlook the crucial role of audio in real-world perception. While several audio-visual benchmarks have been proposed, they face significant limitations. AV-Odyssey Bench [19] and OmniBench [31] focus on static images, Music-AVQA [29] and AVQA [70] are domain-specific with monotonous questions, and LongVALE [18] limits its assessment to captioning capabilities alone. Given that existing benchmarks fail to provide a comprehensive evaluation of MLLMs’ real-world understanding capabilities, we introduce WorldSense to address this critical gap in the field.
3 WorldSense
In this section, we aim to detail the construction of WorldSense, including data collection procedure, annotation pipeline, and statistics. Unlike existing benchmarks that assess modalities in isolation, WorldSense evaluates the MLLMs’ ability to perceive, understand, and reason about real-world scenarios through the integration of omni-modal information. As shown in Figure 1, all the multiple-choice questions are carefully crafted, to make sure the questions can only be answered through the comprehensive analysis of text, vision, and audio.
3.1 Design Principle
As for multi-modal evaluation, we base on the audio-visual synchronized videos, which capture temporal events, motion patterns, and audio-visual correlations. To curate the benchmark, we adhere to the following three principles, to ensure a rigorous and comprehensive evaluation.
Benchmarks Modality #Videos Len.(s) #QA Pairs Anno. QA Tokens Sub. Tokens Multi-task Open-domain Sub./Aud. A-V Correlations MSRVTT-QA [68] V 2,990 15.2 72,821 A 8.4 ✗ ✗ ✓ ✗ ✗ ActivityNet-QA [72] V 800 111.4 8,000 M 10.2 ✗ ✗ ✗ ✗ ✗ MVBench [30] V 3,641 16.0 4,000 A 27.3 ✗ ✓ ✓ ✗ ✗ Video-Bench [47] V 5,917 56.0 17,036 A&M 21.3 ✗ ✓ ✓ ✗ ✗ EgoSchema [44] V 5,063 180.0 5,063 A&M 126.8 ✗ ✓ ✗ ✗ ✗ Video-MME [17] V 900 1017.9 2,700 M 35.7 3086.5 ✓ ✓ ✓ ✗ MMBench-Video [13] V 609 165.4 1,998 M 19.3 ✗ ✓ ✓ ✗ ✗ AVQA [70] A+V 57,000 10 57,335 M 14.2 ✗ ✗ ✓ ✓ ✓ Music-AVQA [29] A+V 9,288 60 45,867 M 8.6 ✗ ✗ ✗ ✓ ✓ OmniBench [31] A+I ✗ ✗ 1,142 M 37.8 ✗ ✓ ✓ ✓ ✓ AV-Odyssey [19] A+I ✗ ✗ 4,555 M 19.5 ✗ ✓ ✓ ✓ ✓ LongVALE [18] A+V 8,400 235 ✗ A&M ✗ ✗ ✗ ✓ ✓ ✓ WorldSense A+V 1,662 141.1 3,172 M 37.2 986.2 ✓ ✓ ✓ ✓
Comprehensive Domain. To ensure a thorough evaluation of MLLMs’ real-world understanding, we developed a systematic taxonomy that covers diverse domains and scenarios. This process started with primary categories reflecting core aspects of human experience, which were further subdivided into 67 subcategories to capture specific contexts. This hierarchical structure ensures that our video collection spans a broad range of real-world experiences, providing an ecologically valid basis for assessing multimodal understanding.
Diverse Acoustic Signals. In real-world scenarios, audio signals can be principally categorized into three fundamental types: speech, event and music. Our benchmark incorporates all three to ensure comprehensive coverage, enabling MLLMs to process and understand a wide spectrum of acoustic information—from semantic speech to abstract music and environmental sounds.
Multilevel Assessment. To evaluate the MLLMs’ perceptual and cognitive capabilities, we designed a multi-scale assessment at three levels: recognition (basic audiovisual element detection), understanding (comprehension of multimodal relationships), and reasoning (advanced cognitive tasks like causal inference and abstract thinking). We developed 26 tasks across to assess different aspects of multimodal comprehension, with a focus on integrating audio-visual information at each level.
3.2 Data Collection & Curation
We primarily source our video content from FineVideo [15], a large-scale dataset of high-quality YouTube videos with strong audio-visual correlations across diverse real-world scenarios. To enhance the coverage of musical content, we additionally incorporate videos from MusicAVQA [29].
Our data collection process involves a systematic filtering pipeline to ensure high-quality videos with rich visual-audio semantics and temporal dynamics, as illustrated in Figure 3(a). The process follows three key steps: (i) filtering videos based on predefined taxonomic categories described in Section 3.1 for comprehensive coverage; (ii) utilizing pre-computed metrics, including audio-visual correlation and dynamic content scores, to identify significant clips from an initial pool of approximately 8,000 videos; and (iii) conducting human expert review to assess video quality and real-world relevance. This rigorous process ultimately yields 1,662 high-quality video segments featuring strong audio-visual correlations across diverse real-world contexts.
3.3 Annotation Protocol
QA Annotation. A team of 80 professional annotators is engaged in creating high-quality multiple-choice QA pairs. For each video clip, annotators conduct thorough reviews of both visual and auditory content to ensure comprehensive understanding. They then generate questions and corresponding answer options that specifically require the integration of both visual and audio information for correct responses, thereby enabling effective evaluation of MLLMs’ multimodal understanding capabilities.
Quality Control. To ensure the quality of question-answer pairs, we implement a rigorous quality control process that combines human expertise with automated verification, as illustrated in Figure 3(b). Professional quality control experts evaluate each QA pair based on three essential criteria: (i) linguistic clarity and coherence, (ii) necessity of both visual and audio information for correct answers, and (iii) appropriateness of the question’s difficulty. Questions that fail to meet these standards are returned for revision.
We additionally employ MLLMs for automated verification. Vision-language models like Qwen2-VL[61] verify that questions require multiple modalities for correct answers. Furthermore, multimodal MLLMs capable of processing video, audio, and text, such as Video-LLaMA2[7] and OneLLM [21] are used to assess question difficulty, with questions answered correctly by all models being flagged for manual revision as too simple.
This dual-verification system, combining expert review and automated testing, ensures that all questions in our benchmark are of high-quality and well-formulated, that requires multi-modal comprehension, and present significant challenges for the models.
Methods LLM Size Tech & Science Culture & Politics Daily Life Film & TV Performance Games Sports Music Avg Open-Source Video-Audio MLLMs Unified-IO-2 L [42] 1B 19.3 22.8 23.1 25.6 25.8 24.1 22.9 25.3 23.3 Unified-IO-2 XL [42] 3B 26.5 24.4 22.5 23.5 24.7 28.0 25.7 24.2 24.7 Unified-IO-2 XXL [42] 7B 27.1 31.7 23.9 23.7 25.5 23.7 25.7 27.3 25.9 OneLLM [21] 7B 26.7 25.1 19.0 22.7 27.0 23.7 22.4 19.8 22.8 VideoLLaMA2 [7] 7B 29.4 25.4 21.8 24.5 26.2 24.6 25.5 27.1 25.4 Open-Source Video MLLMs Video-LLaVA [32] 7B 23.6 20.8 19.1 17.3 23.6 17.2 20.8 20.1 20.3 LLaMA3.2 [20] 7B 27.5 25.7 28.9 25.9 27.7 21.1 29.0 26.8 27.1 Qwen2-VL [62] 7B 33.5 29.0 28.4 33.6 30.3 32.3 34.7 38.5 32.4 mPLUG-Owl3 [71] 7B 37.5 31.4 31.0 34.1 33.3 33.2 32.1 30.5 32.9 LLaVA-OneVision [26] 7B 38.9 38.9 36.3 37.6 37.8 37.9 36.3 39.1 37.7 InternVL2.5 [5] 8B 43.7 40.9 34.6 39.7 37.8 36.2 39.4 41.1 39.1 LLaVA-Video [78] 7B 41.6 38.6 40.6 42.1 40.4 39.7 37.0 40.9 40.2 Proprietary MLLMs Claude 3.5 Sonnet [2] - 43.7 31.7 30.6 36.5 30.7 31.9 36.6 33.9 34.8 GPT 4o [24] - 48.0 44.0 38.3 43.5 41.9 41.2 42.6 42.7 42.6 Gemini 1.5 Pro [58] - 53.7 47.2 50.3 50.4 52.4 46.8 40.2 42.0 48.0
3.4 Dataset Statistics
As summarized in Table 1, the WorldSense dataset contains 1,662 video clips with synchronized audio, distributed across 8 primary categories and 67 subcategories. The average duration is 141.1 seconds, with lengths ranging from 30 seconds to over 10 minutes, capturing a wide variety of events and activities. In total, WorldSense includes 3,173 multiple-choice questions, covering three levels.
WorldSense encompasses diverse audio components including speech, environmental sounds, and music. Unlike existing benchmarks that use static images (e.g., AV-Odyssey Bench [19], OmniBench [31]) or feature weak audio-visual correlations (e.g., Video-MME [17]), WorldSense is the first benchmark designed to evaluate MLLMs’ real-world multimodal understanding. It distinguishes itself through: (i) open-domain videos with multi-task evaluation, (ii) original audio-visual content with complete transcriptions, and (iii) carefully crafted questions requiring true audio-visual integration, establishing a more comprehensive benchmark for real-world multimodal understanding assessment.
3.5 Evaluation Paradigm
In our evaluation framework, each test instance consists of a video clip with synchronized audio and a multiple-choice question. Models must process these multi-modal inputs and select the correct answer from several options. Performance is measured by accuracy, comparing the model’s selection to the ground-truth answers. A model’s success is determined by its ability to accurately align with the correct answer.
To rigorously assess the necessity of multimodal integration in real-world understanding, we conduct ablation studies across various modality configurations. This approach not only evaluates overall model performance but also quantifies the models’ reliance on individual modalities, highlighting the critical role of multimodal collaboration in real-world comprehension tasks.
4 Experiments and Findings
This section presents a comprehensive evaluation of both open-source and proprietary MLLMs on the WorldSense benchmark. We first delineate our experimental methodology and evaluation protocols, followed by a comprehensive analysis of quantitative results. Furthermore, we conduct detailed investigations into the important factors that affect performance, providing insights that illuminate potential directions for multi-modal understanding.
4.1 Settings
To comprehensively assess the multi-modal understanding ability, we evaluate three types of MLLMs: (1) open-source audio-visual models, such as Unified-IO-2 [42], OneLLM [21], and VideoLLaMA2 [7]; (ii) open-source MLLMs, such as Qwen2-VL [62], LLaVA-OneVision [26], InternVL2.5 [5], LLaVA-Video [78], and so on; (iii) proprietary MLLMs, such as Claude 3.5 Sonnet [2], GPT 4o [24], and Gemini 1.5 Pro [58]. For all evaluations, we strictly adhere to each model’s official implementation guidelines and the recommended pre-processing procedures. Video frame extraction follows the official configurations specified by corresponding MLLMs, while proprietary models are evaluated according to their API specifications and recommended input formats. Model performance is assessed through direct comparison between model outputs and ground-truth answers.
4.2 Results on WorldSense
Main Results. We present the comprehensive evaluation results on WorldSense in Table 2. Our findings reveal several significant insights regarding the current state of multi-modal models in real-world understanding.
First, current open-source video models are limited in their performance as they process only visual information. This restriction highlights a significant gap in their ability to perform complex, multi-modal understanding tasks, as evidenced by their maximum performance score of only . The results underscore the inadequacies of relying solely on visual processing, emphasizing the need to integrate audio inputs for a more comprehensive and accurate understanding in practical applications.
Methods Modality Tech & Science Culture & Politics Daily Life Film & TV Performance Games Sports Music Avg Unified-IO-2 L [42] Audio 23.0 25.4 24.2 26.7 27.7 23.7 25.0 27.1 25.2 + Caption 21.5 21.1 20.7 17.1 19.9 19.0 22.9 23.7 20.9-4.3 + Video 19.3 22.8 23.1 25.6 25.8 24.1 22.9 25.3 23.3-1.9 Unified-IO-2 XL [42] Audio 21.7 22.4 22.4 22.1 24.7 25.0 25.9 24.7 23.4 + Caption 19.9 19.8 20.8 19.2 20.2 15.9 21.7 25.5 20.7-2.7 + Video 26.5 24.4 22.5 23.5 24.7 28.0 25.7 24.2 24.7+1.3 Unified-IO-2 XXL [42] Audio 27.5 28.7 23.9 23.2 25.8 21.1 26.2 30.2 25.9 + Caption 24.0 26.7 23.0 18.9 18.7 20.7 25.9 29.4 23.7-2.2 + Video 27.1 31.7 23.9 23.7 25.5 23.7 25.7 27.3 25.9+0.0 OneLLM [21] Audio 25.7 26.1 19.3 21.9 25.8 25.9 21.5 22.4 23.0 + Caption 29.6 29.0 25.9 29.1 33.0 26.7 29.2 28.6 28.6+5.6 + Video 26.7 25.1 19.0 22.7 27.0 23.7 22.4 19.8 22.8-0.2 VideoLLaMA2 [7] Audio 23.8 23.4 21.3 22.4 24.7 19.8 27.1 27.9 23.8 + Caption 30.0 30.0 25.6 29.9 28.5 25.0 29.7 29.9 28.5+4.7 + Video 29.4 25.4 21.8 24.5 26.2 24.6 25.5 27.1 25.4+1.6 Gemini 1.5 Pro [58] Audio 40.2 42.9 35.8 33.3 33.0 31.0 33.3 24.7 34.6 + Caption 49.5 52.1 41.8 42.9 46.4 41.8 39.6 36.7 43.6+9.0 + Video 53.7 47.2 50.3 50.4 52.4 46.8 40.2 42.0 48.0+13.4
Second and surprisingly, existing open-source audio-visual MLLMs perform even worse, achieving accuracy rates comparable to random guessing and notably below video-only MLLMs. This counter-intuitive finding reveals that despite having access to both modalities, these models struggle with effective audio-visual integration, suggesting that multimodal processing capability alone does not guarantee better performance without sophisticated integration mechanisms.
Third, among proprietary MLLMs, vision-only models GPT-4o and Claude 3.5 Sonnet perform similarly to the best open-source video MLLMs. Gemini 1.5 Pro, capable of processing both audio and visual information, achieves the highest accuracy of . However, this performance still falls considerably short of requirements for reliable real-world applications, indicating substantial room for improvement.
These comprehensive results illuminate several critical insights: (i) the fundamental importance of audio-visual collaborative understanding in real-world scenarios; (ii) the current significant gap in models’ capabilities for effective multimodal integration, and (iii) the need for more sophisticated approaches to combining and reasoning about multiple modalities. These findings point to crucial directions for future research and development in MLLMs.
Breakdown Results. We conduct a fine-grained analysis of model performance across different audio types and task categories, as shown in Figure 4, revealing important insights into the limitations of existing multimodal models.
First, models consistently underperform on audio-related tasks (e.g., audio recognition, audio counting) compared to other task types, demonstrating significant challenges in audio understanding. Second, spatial reasoning and counting tasks present notable difficulties for current models, a pattern consistently observed across multiple benchmarks. Third, emotion-related tasks prove particularly challenging, likely due to their requirement for integrating subtle and complex multimodal cues, including facial expressions, vocal tones, and contextual speech content. This underperformance in emotional understanding suggests a significant gap in current MLLMs’ training data and capabilities, highlighting an important area for future development.
Additionally, performance varies across audio types. While Gemini 1.5 Pro performs best overall, it shows notably lower accuracy on event-related questions compared to speech or music tasks, possibly due to the complex nature of environmental sounds. Other models also exhibit inconsistent performance across audio types, underscoring a general limitation for existing models for audio understanding.
4.3 Roadmap Towards Real-world Understanding
Given the substantial performance gap revealed in above evaluation, we conduct an in-depth investigation into potential approaches to enhance the MLLMs’ performance.
Vision Information. We investigate the impact of visual information through different input configurations: audio-only, audio with video captions, and audio with video frames. As shown in Table 3, visual information generally improves performance, with Gemini 1.5 Pro’s accuracy increasing from (audio-only) to (+video). However, impact varies across models, with UnifiedIO2 showing inconsistent gains and even degradation with captions.
These findings suggest two important insights: (1) visual information is crucial for enhancing multi-modal understanding when properly integrated, and (2) current models’ ability to effectively utilize visual information remains limited.
Audio Information. We examine the impact of audio information through three configurations: video-only, video with subtitles, and video with original audio.
The results in Table 5 reveal intriguing patterns in how different forms of audio information influence model performance. For Gemini 1.5 Pro, accuracy increases from (video-only) to with subtitles, and further to with original audio. OneLLM shows similar improvements. These results demonstrate that both subtitles and acoustic features (including tone, emotion, and environmental sounds) contribute valuable information for multimodal understanding, beyond what subtitles alone can capture, highlighting the importance of complete acoustic cues for multimodal understanding.
Interestingly, UnifiedIO2 demonstrates performance degradation when integrating either subtitles or audio, with subtitles causing a notable accuracy decline, suggesting difficulties in multimodal processing. Conversely, Video-LLaMA2 improves with both modalities but performs better with subtitles than original audio, indicating a stronger reliance on textual rather than complex acoustic information. We further evaluate video-only MLLMs by providing transcribed subtitles, as shown in Table 5. Nearly all models show significant improvements with subtitle integration, reinforcing the importance of audio information. However, the performance gain is less pronounced in music-related questions, as subtitles cannot effectively capture inherent acoustic features such as melody, rhythm, and harmony.
These evaluations highlight several critical findings: (i) original audio contains rich information beyond what subtitles can capture, particularly for music; (ii) current models show significant limitations in multimodal processing. These insights suggest important directions for improving MLLMs’ ability to integrate acoustic and textual information for comprehensive scene understanding.
Methods Speech Event Music Overall Video + Subtitle + Audio Video + Subtitle + Audio Video + Subtitle + Audio Video + Subtitle + Audio Unified-IO-2 L [42] 26.8 13.9-12.9 23.1-3.1 26.9 13.5-13.4 23.2-3.7 26.3 15.0-11.3 24.5-1.8 26.6 14.8-11.8 23.3-3.3 Unified-IO-2 XL [42] 25.0 13.0-12.0 24.6-0.4 24.8 12.3-12.5 25.0+0.2 26.7 15.9-10.8 24.7-2.0 25.3 14.1-11.2 24.7-0.6 Unified-IO-2 XXL [42] 27.0 15.6-11.4 25.7-1.3 26.2 14.2-12.0 25.1-1.1 28.4 19.1-9.3 25.9-2.5 27.2 17.2-10.0 25.9-1.3 OneLLM [21] 12.5 19.6+7.1 23.2+10.7 12.4 19.3+6.9 22.3+9.9 12.4 19.0+6.6 22.5+10.1 12.6 19.6+7.0 22.8+10.2 VideoLLaMA2 [7] 17.1 25.5+8.4 25.2+8.1 16.1 24.9+8.8 24.5+8.4 17.7 27.0+9.3 25.6+7.9 17.4 26.1+8.7 25.4+8.0 Gemini 1.5 Pro [58] 34.3 39.6+5.3 49.2+14.9 33.0 38.9+5.9 48.7+15.7 35.4 39.2+3.8 49.8+14.4 34.4 39.3+4.9 48.0+13.6
Methods Speech Event Music Overall Video + Subtitle Video + Subtitle Video + Subtitle Video + Subtitle Video-LLaVA [32] 20.3 15.4-4.9 19.8 14.4 -5.4 19.5 16.4-3.1 20.3 16.0-4.3 LLaMA3.2 [20] 27.1 29.3+2.2 27.6 29.6+2.0 25.9 28.1+2.2 27.1 28.8+1.7 Qwen2-VL [62] 31.8 41.1+9.3 30.9 39.4+8.5 34.2 41.8+7.6 32.4 41.2+8.8 mPLUG-Owl3 [71] 33.0 39.2+6.2 32.3 38.3+6.0 34.6 39.2+4.6 32.9 38.7+5.8 LLaVA-OneVision [26] 37.7 44.0+6.3 36.3 42.7+6.4 39.7 45.7+6.0 37.7 43.9+6.2 InternVL2.5 [5] 39.0 48.3+9.3 38.6 47.9+9.3 39.2 47.1+7.9 39.1 47.8+8.7 LLaVA-Video [78] 40.5 45.9+5.4 38.9 44.6+5.7 42.3 47.7+5.4 40.2 45.6+5.4 GPT 4o [24] 42.8 51.1+8.3 40.9 50.2+9.3 43.6 49.9+6.3 42.6 50.1+7.5
Video Frames. We investigate the effect of temporal sampling density by varying input frame numbers for video-only MLLMs. As shown in Figure 5, most models demonstrate significant performance improvements with increased frame density, with LLaMA-3.2 being a notable exception. These gains likely result from better capture of fine-grained temporal dynamics and subtle visual changes, emphasizing the importance of dense temporal sampling.
5 Conclusion
In this paper, we introduce WorldSense, the first benchmark designed to evaluate MLLMs’ omnimodal understanding in real-world scenarios. Distinguished by its emphasis on joint omnimodal comprehension across diverse real-world contexts, WorldSense encompasses rich video categories and carefully curated question-answer pairs that necessitate the integration of visual and acoustic information. Through extensive experiments, we expose significant limitations in current MLLMs’ ability to process and coherently integrate omnimodal information. Our analysis demonstrates the importance of omnimodal collaboration in real-world understanding. We hope that WorldSense can serve as a foundational benchmark for advancing human-like omnimodal understanding capabilities.
Appendix A Broader Impacts
Our work on WorldSense has several potential positive impacts on society and AI development, while also presenting certain risks that warrant careful consideration. WorldSense contributes to advancing MLLMs’ ability to understand and interact with the real world through multiple modalities. This progress could benefit various applications, including assistive technologies, educational tools, human-AI interaction systems, safety systems, and so on. We also acknowledge potential risks and challenges. The development of more capable AI systems might raise privacy concerns. Advanced multimodal understanding capabilities could potentially be misused for surveillance or monitoring purposes. We believe that open discussion of these impacts is crucial for responsible development of MLLMs.
Appendix B Ethics Statement
Our research on WorldSense adheres to strict ethical principles and guidelines. We acknowledge several important ethical considerations:
Data Collection and Privacy. All video content in WorldSense has been collected from publicly available sources with appropriate licensing agreements. We have con ducted thorough reviews and implemented comprehensive data processing procedures to ensure privacy protection, including the removal of any personally identifiable information.
Potential Biases. While acknowledging that inherent biases may exist in any dataset, we have undertaken systematic efforts to ensure diverse representation across our video content and question-answer pairs, encompassing various domains, cultures, and contexts. Nevertheless, we recognize that completely eliminating bias remains a significant challenge, and users should carefully consider these potential limitations when utilizing our dataset.
Intended Use. WorldSense is specifically designed to advance research in omnimodal real-world understanding. While we actively encourage the use of this benchmark for academic and research purposes, we strongly caution against any applications that could potentially result in harmful or discriminatory outcomes. Users are expected to adhere to ethical guidelines and responsible practices.
Appendix C License
The WorldSense dataset is released under the CC BY-NC-SA 4.0 License. Authors bear all responsibility in case of violation of rights and confirmation of the data license.

### v2

@@
 Abstract
-In this paper, we introduce WorldSense, the first benchmark to assess the multi-modal video understanding, that simultaneously encompasses visual, audio, and text inputs.
+We introduce WorldSense, the first benchmark to assess the multi-modal video understanding, that simultaneously encompasses visual, audio, and text inputs.
 In contrast to existing benchmarks, our WorldSense has several features: (i) collaboration of omni-modality, we design the evaluation tasks to feature a strong coupling of audio and video, requiring models to effectively utilize the synergistic perception of omni-modality; (ii) diversity of videos and tasks, WorldSense encompasses a diverse collection of 1,662 audio-visual synchronised videos, systematically categorized into 8 primary domains and 67 fine-grained subcategories to cover the broad scenarios, and 3,172 multi-choice QA pairs across 26 distinct tasks to enable the comprehensive evaluation; (iii) high-quality annotations, all the QA pairs are manually labeled by 80 expert annotators with multiple rounds of correction to ensure quality.
@@
 The experimental results indicate that existing models face significant challenges in understanding real-world scenarios ( best accuracy).
+By analyzing the limitations of current models, we aim to provide valuable insight to guide development of real-world understanding.
 We hope our WorldSense can provide a platform for evaluating the ability in constructing and understanding coherent contexts from omni-modality.
@@
 For example, when driving a car, a human driver integrates visual information (e.g., recognizing road signs, traffic lights, and obstacles), auditory cues (e.g., hearing the honking of another car or a siren approaching from behind), and tactile feedback (e.g., the feel of the steering wheel, the vibrations of the road, or the responsiveness of the brakes) to make real-time decisions and ensure safe navigation.
-This seamless multimodal integration enables intelligent agents to process complex, dynamic environments and respond to subtle cues—an ability that is essential for both human perception and the development of embodied agents designed to interact naturally in the world.
+This seamless multimodal integration enables intelligent agents to process complex, dynamic environments and respond to subtle cues—an ability that is essential for both human perception and development of embodied agents designed to interact naturally.
 In the recent literature, the development of Multi-modal Large Language Models (MLLMs) [51, 25, 50, 59, 61, 79, 44, 12] have led to remarkable progress on a series of tasks, for example, classification [37], captioning [1, 10, 36], question-answering [58, 52, 41], OCR [47, 80], segmentation [26, 69, 23], autonomous driving [48, 54, 4] and more.
 However, multi-modal analysis primarily focuses on visual-language information, leaving out crucial modalities like audio, which results in an incomplete evaluation of their multimodal capabilities.
-While some benchmarks have begun incorporating both visual and audio modalities, they still exhibit several limitations.
+While some benchmarks have started incorporating both visual and audio modalities, they still exhibit several limitations.
 For example, OmniBench [32] and AV-Odyssey Bench [20] mainly emphasize image evaluation, whereas other benchmarks [19, 30, 74] either restrict to captioning tasks or are limited to simple scenarios, or suffer from low-quality, monotonous questioning patterns.
-In this paper, we introduce WorldSense, the first benchmark designed to comprehensively evaluate MLLMs’ ability to perceive, understand, and reason through the integration of omni-modal information in real-world contexts.
-WorldSense is defined by three key features: (i) Collaboration of Omni-Modality: At the core of WorldSense is the emphasis on the interplay between audio and visual information, as illustrated in Figure 1.
-Each question is designed to require both modalities for a correct answer; the absence of either modality leads to an incorrect response.
-Such design rigorously tests a model’s ability to simultaneously process multiple sensory inputs for accurate understanding; (ii) Diversity of Videos and Tasks: WorldSense encompasses 1,662 audio-visual synchronized videos across 8 domains and 67 fine-grained subcategories.
-It includes 3,172 multiple-choice questions spanning 26 distinct cognitive tasks, from basic perception to complex reasoning.
-This diversity enables a systematic and comprehensive evaluation of MLLMs’ multimodal understanding capabilities; (iii) High-Quality Annotations: To ensure reliability, all QA pairs in WorldSense are manually labeled by 80 expert annotators and undergo multiple rounds of validation through both human review and automatic MLLM verification.
-This rigorous process ensures the benchmark’s accuracy and reliability.
-By integrating these features, WorldSense sets a new standard for evaluating MLLMs’ ability to handle real-world complexity, bridging the gap between artificial intelligence and human-like multimodal understanding.
-We conduct comprehensive experiments on WorldSense to evaluate a range of existing MLLMs, including open-source video models, video-audio models, and proprietary models.
-Our results highlight significant limitations in the current models’ ability to understand and reason through omni-modal information in real-world contexts.
-First, while open-source video-audio models can process both video and audio inputs, they achieve only around 25% accuracy on our benchmark, comparable to random guessing.
-In contrast, proprietary models, such as Gemini 1.5 Pro, which processes both video and audio, achieve the highest accuracy at 48%.
-However, when provided with only a single modality (either video or audio), Gemini 1.5 Pro’s performance drops significantly by approximately 15%, underscoring the crucial role of omni-modal collaboration in real-world tasks.
-These findings emphasize the strong coupling of modalities in WorldSense and reveal a substantial gap in current MLLMs’ real-world understanding capabilities.
-Additionally, we further conduct detailed ablation studies on how different modalities impact model performance.
-Our analysis shows that visual information plays significant for understanding, while adding audio—whether as raw signals or transcribed subtitles—also improves results.
-However, raw audio inputs yield superior gains, emphasizing the importance of paralinguistic features such as prosody, intonation, and acoustic context, which are often lost in text transcription.
-Additionally, increasing the density of video frame sampling, which provides richer temporal information, consistently improves performance.
-These findings underscore the complementary nature of visual and auditory modalities and highlight the necessity of integrating both for robust real-world understanding.
-To summarise, in this paper, we have made three contributions: (i) we introduce, WorldSense, the first benchmark specifically designed to evaluate MLLMs’ omni-modal understanding capabilities in real-world scenarios.
-WorldSense is characterized by the integration of omni-modality, diverse video categories, and high-quality QA pairs; (ii) We have conducted extensive experiments to evaluate current MLLMs’ capabilities in real-world omni-modal understanding.
-Our results demonstrate that open-source video-audio MLLMs perform only marginally better than random guessing, while even the best-performing proprietary models achieve only accuracy, revealing significant room for accurate omni-modal understanding; (iii) Through detailed analysis, we investigate the crucial factors affecting real-world omni-modal understanding.
-Our findings demonstrate that acoustic information, visual cues, and dense temporal sampling all substantially impact model performance.
-These finds provide valuable guidance for the future development of real-world omni-modal understanding.
+This paper presents WorldSense, the first comprehensive benchmark designed to evaluate Multimodal Large Language Models (MLLMs) in perceiving, understanding, and reasoning with omni-modal information in real-world settings.
+The benchmark is defined by three key features: (i) Omni-modal integration.
+THe benchmark emphasizes the joint processing of audio and visual modalities, as illustrated in Figure 1.
+Each question requires both modalities for accurate response—removing either results in failure—enabling rigorous assessment of a model’s capacity for integrated sensory understanding.
+(ii) Diverse videos and task coverage.
+The benchmark includes 1,662 synchronized audio-visual videos spanning 8 domains and 67 fine-grained subcategories.
+It features 3,172 multiple-choice questions across 26 cognitive tasks, ranging from basic perception to high-level reasoning.
+This diversity supports systematic evaluation of multimodal comprehension across a broad task spectrum.
+(iii) High-quality annotations.
+All question-answer pairs are curated by 80 expert annotators and undergo multiple validation rounds, including human review and automated MLLM verification.
+This ensures annotation accuracy and benchmark reliability.
+Through these methodological advancements, WorldSense sets a new standard for evaluating MLLMs in real-world multimodal reasoning, advancing the field toward more human-like understanding.
+We conduct extensive evaluations for a broad spectrum of MLLMs, including open-source video models, video-audio models, and proprietary systems.
+Results reveal significant limitations in current models’ ability to reason over omni-modal inputs in real-world contexts.
+Specifically, open-source video-audio models, despite processing both modalities, achieve only 25% accuracy—comparable to random guessing.
+In contrast, proprietary models such as Gemini 1.5 Pro reach up to 48% accuracy.
+However, when restricted to a single modality (audio or video), Gemini’s performance drops by 15%, highlighting the critical role of integrated modality processing.
+We further conduct ablation studies to dissect modality contributions.
+Visual inputs are essential, while audio—especially raw signals—yields additional gains over text transcriptions, due to preserved paralinguistic cues, e.g., prosody, intonation, acoustic context.
+These findings affirm the complementary nature of audio-visual information and the necessity of their joint modeling for robust real-world understanding.
+Failure case analysis reveals persistent limitations in current MLLMs, motivating future directions for improving multimodal reasoning.
+To summarize, we have made the following contributions: (i) we present WorldSense, the first benchmark tailored for evaluating MLLMs’ ability on omni-modal video understanding, characterized by integrated audio-visual inputs, diverse content, and high-quality question-answering annotations; (ii) we have conducted extensive evaluation of existing MLLMs, showing that most open-source models perform near chance, and even the best proprietary model achieves only 48% accuracy—exposing a significant gap in real-world omni-modal reasoning; (iii) through ablation and failure analysis, we identify the key factors influencing performance, including raw audio and visual cues, and provide actionable insights to guide future omni-modal understanding design.
 2 Related Work
@@
 Current Large Language Models (LLMs) are capable of processing multimodal information, including visual, text, and audio.
-Early works, such as [75, 35, 80, 11, 64, 51], successfully combine vision and text modalities.
-Subsequent research extends to temporal understanding [66, 24, 58, 39, 61, 65, 26, 14, 69, 74, 60, 6, 41, 34], while parallel efforts [55, 9, 8] focus on audio processing.
+Early works, such as [79, 36, 84, 11, 66, 53, 45, 12], successfully combine vision and text modalities.
+Subsequent research extends to temporal understanding [68, 25, 60, 40, 63, 67, 27, 14, 72, 78, 62, 6, 42, 35, 73], while parallel efforts [57, 9, 8] focus on audio processing.
 Recently, researchers shift attention to models [7, 56, 60, 43, 61] capable of simultaneously processing text, vision, and audio inputs.
@@
 3 WorldSense
-In this section, we aim to detail the construction of WorldSense, including data collection procedure, annotation pipeline, and statistics.
-Unlike existing benchmarks that assess modalities in isolation, WorldSense evaluates the MLLMs’ ability to perceive, understand, and reason about real-world scenarios through the integration of omni-modal information.
-As shown in Figure 1, all the multiple-choice questions are carefully crafted, to make sure the questions can only be answered through the comprehensive analysis of text, vision, and audio.
+In this section, we first introduce the design principles in Section 3.1, followed by a description of the data collection (Section 3.2) and annotation processes ( 3.3).
+We then compare statics of WorldSense with previous benchmarks in Section 3.4, and finally present our evaluation methodology 3.5.
 3.1 Design Principle
 As for multi-modal evaluation, we base on the audio-visual synchronized videos, which capture temporal events, motion patterns, and audio-visual correlations.
-To curate the benchmark, we adhere to the following three principles, to ensure a rigorous and comprehensive evaluation.
-Benchmarks Modality #Videos Len.(s) #QA Pairs Anno.
-QA Tokens Sub.
-Tokens Multi-task Open-domain Sub./Aud.
-A-V Correlations MSRVTT-QA [68] V 2,990 15.2 72,821 A 8.4 ✗ ✗ ✓ ✗ ✗ ActivityNet-QA [72] V 800 111.4 8,000 M 10.2 ✗ ✗ ✗ ✗ ✗ MVBench [30] V 3,641 16.0 4,000 A 27.3 ✗ ✓ ✓ ✗ ✗ Video-Bench [47] V 5,917 56.0 17,036 A&M 21.3 ✗ ✓ ✓ ✗ ✗ EgoSchema [44] V 5,063 180.0 5,063 A&M 126.8 ✗ ✓ ✗ ✗ ✗ Video-MME [17] V 900 1017.9 2,700 M 35.7 3086.5 ✓ ✓ ✓ ✗ MMBench-Video [13] V 609 165.4 1,998 M 19.3 ✗ ✓ ✓ ✗ ✗ AVQA [70] A+V 57,000 10 57,335 M 14.2 ✗ ✗ ✓ ✓ ✓ Music-AVQA [29] A+V 9,288 60 45,867 M 8.6 ✗ ✗ ✗ ✓ ✓ OmniBench [31] A+I ✗ ✗ 1,142 M 37.8 ✗ ✓ ✓ ✓ ✓ AV-Odyssey [19] A+I ✗ ✗ 4,555 M 19.5 ✗ ✓ ✓ ✓ ✓ LongVALE [18] A+V 8,400 235 ✗ A&M ✗ ✗ ✗ ✓ ✓ ✓ WorldSense A+V 1,662 141.1 3,172 M 37.2 986.2 ✓ ✓ ✓ ✓
-Comprehensive Domain.
-To ensure a thorough evaluation of MLLMs’ real-world understanding, we developed a systematic taxonomy that covers diverse domains and scenarios.
-This process started with primary categories reflecting core aspects of human experience, which were further subdivided into 67 subcategories to capture specific contexts.
-This hierarchical structure ensures that our video collection spans a broad range of real-world experiences, providing an ecologically valid basis for assessing multimodal understanding.
-Diverse Acoustic Signals.
-In real-world scenarios, audio signals can be principally categorized into three fundamental types: speech, event and music.
-Our benchmark incorporates all three to ensure comprehensive coverage, enabling MLLMs to process and understand a wide spectrum of acoustic information—from semantic speech to abstract music and environmental sounds.
-Multilevel Assessment.
-To evaluate the MLLMs’ perceptual and cognitive capabilities, we designed a multi-scale assessment at three levels: recognition (basic audiovisual element detection), understanding (comprehension of multimodal relationships), and reasoning (advanced cognitive tasks like causal inference and abstract thinking).
-We developed 26 tasks across to assess different aspects of multimodal comprehension, with a focus on integrating audio-visual information at each level.
+To curate the benchmark, we adhere to the following three principles, to ensure rigorous and comprehensive evaluations for MLLMs.
+Comprehensive Domain Coverage.
+To capture the diversity of real-world scenarios, we construct a hierarchical taxonomy starting from broad human-centric domains, refined into 67 fine-grained subcategories.
+This structure ensures wide ecological coverage, enabling robust assessment of multi-modal understanding across varied contexts.
+Diverse Acoustic Modalities.
+Real-world audio can be broadly classified into speech, environmental events, and music.
+The benchmark includes all three types, enabling evaluation across a spectrum of acoustic complexity—from linguistic content to non-verbal and abstract auditory cues.
+Multilevel Cognitive Assessment.
+We design a three-tiered evaluation framework targeting: recognition (detection of basic audio-visual elements), understanding (comprehension of multimodal relationships), and reasoning (high-level inference tasks such as causal inference or abstract thinking).
+The benchmark includes 26 tasks aligned with these levels, encouraging holistic evaluation of perceptual and cognitive capabilities in multimodal settings.
 3.2 Data Collection & Curation
-We primarily source our video content from FineVideo [15], a large-scale dataset of high-quality YouTube videos with strong audio-visual correlations across diverse real-world scenarios.
-To enhance the coverage of musical content, we additionally incorporate videos from MusicAVQA [29].
-Our data collection process involves a systematic filtering pipeline to ensure high-quality videos with rich visual-audio semantics and temporal dynamics, as illustrated in Figure 3(a).
-The process follows three key steps: (i) filtering videos based on predefined taxonomic categories described in Section 3.1 for comprehensive coverage; (ii) utilizing pre-computed metrics, including audio-visual correlation and dynamic content scores, to identify significant clips from an initial pool of approximately 8,000 videos; and (iii) conducting human expert review to assess video quality and real-world relevance.
-This rigorous process ultimately yields 1,662 high-quality video segments featuring strong audio-visual correlations across diverse real-world contexts.
+We primarily source our video content from FineVideo [15], a large-scale dataset comprising high-quality YouTube videos that exhibit strong audio-visual correlations across diverse real-world scenarios.
+To enrich the benchmark’s coverage of musical content, we supplement it with selected videos from MusicAVQA [30], ensuring a more balanced representation of auditory modalities.
+Our data collection employs a systematic filtering pipeline to ensure high-quality videos with rich visual-audio semantics and temporal dynamics, following three main steps in Figure 3(a): (i) filtering videos according to predefined taxonomic categories delineated in Section 3.1; (ii) selecting clips based on pre-computed audio-visual correlation and dynamic content metrics from about 8,000 initial videos; and (iii) human expert review for video quality and real-world relevance.
+This rigorous selection and processing results in 1,662 high-quality video segments with strong audio-visual correlations across various real-world scenarios.
 3.3 Annotation Protocol
-QA Annotation.
-A team of 80 professional annotators is engaged in creating high-quality multiple-choice QA pairs.
-For each video clip, annotators conduct thorough reviews of both visual and auditory content to ensure comprehensive understanding.
-They then generate questions and corresponding answer options that specifically require the integration of both visual and audio information for correct responses, thereby enabling effective evaluation of MLLMs’ multimodal understanding capabilities.
+Question-Answering (QA) Annotation.
+A team of 80 professional annotators is engaged in creating high-quality multiple-choice QA pairs for each video by thoroughly reviewing both visual and audio content.
+The questions are designed to require integration of multiple modalities, enabling effective assessment of MLLMs’ multimodal understanding.
 Quality Control.
-To ensure the quality of question-answer pairs, we implement a rigorous quality control process that combines human expertise with automated verification, as illustrated in Figure 3(b).
-Professional quality control experts evaluate each QA pair based on three essential criteria: (i) linguistic clarity and coherence, (ii) necessity of both visual and audio information for correct answers, and (iii) appropriateness of the question’s difficulty.
+To ensure QA quality, we implement a rigorous quality control process combining expert review and automated checks, as illustrated in Figure 3(b).
+Professional quality control experts evaluate each QA pair based on three essential criteria: (i) linguistic clarity and coherence, (ii) multimodal necessity for correct answers, and (iii) appropriate difficulty.
 Questions that fail to meet these standards are returned for revision.
-We additionally employ MLLMs for automated verification.
+We also use MLLMs for automated verification.
 Vision-language models like Qwen2-VL[63] verify that questions require multiple modalities for correct answers.
@@
 This dual-verification system, combining expert review and automated testing, ensures that all questions in our benchmark are of high-quality and well-formulated, that requires multi-modal comprehension, and present significant challenges for the models.
-Methods LLM Size Tech & Science Culture & Politics Daily Life Film & TV Performance Games Sports Music Avg Open-Source Video-Audio MLLMs Unified-IO-2 L [42] 1B 19.3 22.8 23.1 25.6 25.8 24.1 22.9 25.3 23.3 Unified-IO-2 XL [42] 3B 26.5 24.4 22.5 23.5 24.7 28.0 25.7 24.2 24.7 Unified-IO-2 XXL [42] 7B 27.1 31.7 23.9 23.7 25.5 23.7 25.7 27.3 25.9 OneLLM [21] 7B 26.7 25.1 19.0 22.7 27.0 23.7 22.4 19.8 22.8 VideoLLaMA2 [7] 7B 29.4 25.4 21.8 24.5 26.2 24.6 25.5 27.1 25.4 Open-Source Video MLLMs Video-LLaVA [32] 7B 23.6 20.8 19.1 17.3 23.6 17.2 20.8 20.1 20.3 LLaMA3.2 [20] 7B 27.5 25.7 28.9 25.9 27.7 21.1 29.0 26.8 27.1 Qwen2-VL [62] 7B 33.5 29.0 28.4 33.6 30.3 32.3 34.7 38.5 32.4 mPLUG-Owl3 [71] 7B 37.5 31.4 31.0 34.1 33.3 33.2 32.1 30.5 32.9 LLaVA-OneVision [26] 7B 38.9 38.9 36.3 37.6 37.8 37.9 36.3 39.1 37.7 InternVL2.5 [5] 8B 43.7 40.9 34.6 39.7 37.8 36.2 39.4 41.1 39.1 LLaVA-Video [78] 7B 41.6 38.6 40.6 42.1 40.4 39.7 37.0 40.9 40.2 Proprietary MLLMs Claude 3.5 Sonnet [2] - 43.7 31.7 30.6 36.5 30.7 31.9 36.6 33.9 34.8 GPT 4o [24] - 48.0 44.0 38.3 43.5 41.9 41.2 42.6 42.7 42.6 Gemini 1.5 Pro [58] - 53.7 47.2 50.3 50.4 52.4 46.8 40.2 42.0 48.0
 3.4 Dataset Statistics
-As summarized in Table 1, the WorldSense dataset contains 1,662 video clips with synchronized audio, distributed across 8 primary categories and 67 subcategories.
-The average duration is 141.1 seconds, with lengths ranging from 30 seconds to over 10 minutes, capturing a wide variety of events and activities.
-In total, WorldSense includes 3,173 multiple-choice questions, covering three levels.
-WorldSense encompasses diverse audio components including speech, environmental sounds, and music.
-Unlike existing benchmarks that use static images (e.g., AV-Odyssey Bench [19], OmniBench [31]) or feature weak audio-visual correlations (e.g., Video-MME [17]), WorldSense is the first benchmark designed to evaluate MLLMs’ real-world multimodal understanding.
-It distinguishes itself through: (i) open-domain videos with multi-task evaluation, (ii) original audio-visual content with complete transcriptions, and (iii) carefully crafted questions requiring true audio-visual integration, establishing a more comprehensive benchmark for real-world multimodal understanding assessment.
+As summarized in Table 1, our proposed WorldSense benchmark contains 1,662 video clips with synchronized audio across 8 categories and 67 subcategories, averaging 141.1 seconds in length, including 3,173 multiple-choice questions on three cognitive levels.
+Benchmarks Modality #Videos Len.(s) #QA Pairs Anno.
+QA Tokens Sub.
+Tokens Multi task Open domain Sub./ Aud.
+A-V Correlations MSRVTT-QA [70] V 2,990 15.2 72,821 A 8.4 ✗ ✗ ✓ ✗ ✗ ActivityNet-QA [76] V 800 111.4 8,000 M 10.2 ✗ ✗ ✗ ✗ ✗ MVBench [31] V 3,641 16.0 4,000 A 27.3 ✗ ✓ ✓ ✗ ✗ MovieChat [55] V 130 500.0 1,950 M - ✗ ✗ ✓ ✗ ✗ Video-Bench [49] V 5,917 56.0 17,036 A&M 21.3 ✗ ✓ ✓ ✗ ✗ EgoSchema [46] V 5,063 180.0 5,063 A&M 126.8 ✗ ✓ ✗ ✗ ✗ Video-MME [17] V 900 1017.9 2,700 M 35.7 3086.5 ✓ ✓ ✓ ✗ MMBench-Video [13] V 609 165.4 1,998 M 19.3 ✗ ✓ ✓ ✗ ✗ AVQA [74] A+V 57,000 10 57,335 M 14.2 ✗ ✗ ✓ ✓ ✓ Music-AVQA [30] A+V 9,288 60 45,867 M 8.6 ✗ ✗ ✗ ✓ ✓ OmniBench [32] A+I ✗ ✗ 1,142 M 37.8 ✗ ✓ ✓ ✓ ✓ AV-Odyssey [20] A+I ✗ ✗ 4,555 M 19.5 ✗ ✓ ✓ ✓ ✓ LongVALE [19] A+V 8,400 235 ✗ A&M ✗ ✗ ✗ ✓ ✓ ✓ WorldSense A+V 1,662 141.1 3,172 M 37.2 986.2 ✓ ✓ ✓ ✓
+WorldSense features diverse audio types such as speech, environmental sounds, and music.
+Unlike existing benchmarks that use static images (e.g., AV-Odyssey Bench [20], OmniBench [32]) or feature weak audio-visual correlations (e.g., Video-MME [17]), WorldSense is the first to comprehensively evaluate MLLMs’ real-world multimodal understanding.
+It distinguishes itself through: (i) open-domain videos with multi-task evaluation, (ii) original audio-visual content with complete transcriptions, and (iii) carefully crafted questions requiring true audio-visual integration, establishing a comprehensive benchmark for real-world multimodal understanding assessment.
 3.5 Evaluation Paradigm
@@
 A model’s success is determined by its ability to accurately align with the correct answer.
+We employ a matching-based approach to extract answers.
 To rigorously assess the necessity of multimodal integration in real-world understanding, we conduct ablation studies across various modality configurations.
@@
 4 Experiments and Findings
-This section presents a comprehensive evaluation of both open-source and proprietary MLLMs on the WorldSense benchmark.
-We first delineate our experimental methodology and evaluation protocols, followed by a comprehensive analysis of quantitative results.
-Furthermore, we conduct detailed investigations into the important factors that affect performance, providing insights that illuminate potential directions for multi-modal understanding.
 4.1 Settings
@@
 Model performance is assessed through direct comparison between model outputs and ground-truth answers.
+Methods LLM Size Tech & Science Culture & Politics Daily Life Film & TV Perfor- mance Games Sports Music Avg Open-Source Video-Audio MLLMs Unified-IO-2 L [43] 1B 19.3 22.8 23.1 25.6 25.8 24.1 22.9 25.3 23.3 Unified-IO-2 XL [43] 3B 26.5 24.4 22.5 23.5 24.7 28.0 25.7 24.2 24.7 Unified-IO-2 XXL [43] 7B 27.1 31.7 23.9 23.7 25.5 23.7 25.7 27.3 25.9 OneLLM [22] 7B 26.7 25.1 19.0 22.7 27.0 23.7 22.4 19.8 22.8 VideoLLaMA2 [7] 7B 29.4 25.4 21.8 24.5 26.2 24.6 25.5 27.1 25.4 VITA-1.5 [18] 7B 38.2 35.9 34.3 39.8 41.2 32.6 34.7 39.9 36.9 Qwen2.5-Omni [71] 7B 47.8 49.8 43.6 43.8 48.3 39.1 43.5 47.3 45.4 Open-Source Video MLLMs Video-LLaVA [33] 7B 23.6 20.8 19.1 17.3 23.6 17.2 20.8 20.1 20.3 LLaMA3.2 [21] 7B 27.5 25.7 28.9 25.9 27.7 21.1 29.0 26.8 27.1 Qwen2-VL [64] 7B 33.5 29.0 28.4 33.6 30.3 32.3 34.7 38.5 32.4 mPLUG-Owl3 [75] 7B 37.5 31.4 31.0 34.1 33.3 33.2 32.1 30.5 32.9 LLaVA-OneVision [27] 7B 38.9 38.9 36.3 37.6 37.8 37.9 36.3 39.1 37.7 InternVL2.5 [5] 8B 43.7 40.9 34.6 39.7 37.8 36.2 39.4 41.1 39.1 LLaVA-Video [82] 7B 41.6 38.6 40.6 42.1 40.4 39.7 37.0 40.9 40.2 Proprietary MLLMs Claude 3.5 Sonnet [2] - 43.7 31.7 30.6 36.5 30.7 31.9 36.6 33.9 34.8 GPT 4o [25] - 48.0 44.0 38.3 43.5 41.9 41.2 42.6 42.7 42.6 Gemini 1.5 Pro [60] - 53.7 47.2 50.3 50.4 52.4 46.8 40.2 42.0 48.0
 4.2 Results on WorldSense
-Main Results.
-We present the comprehensive evaluation results on WorldSense in Table 2.
-Our findings reveal several significant insights regarding the current state of multi-modal models in real-world understanding.
+Main Results. We present comprehensive evaluations of WorldSense in Table 2. Our analysis reveals several significant insights regarding the current capabilities of MLLMs in real-world understanding.
 First, current open-source video models are limited in their performance as they process only visual information.
@@
 The results underscore the inadequacies of relying solely on visual processing, emphasizing the need to integrate audio inputs for a more comprehensive and accurate understanding in practical applications.
-Methods Modality Tech & Science Culture & Politics Daily Life Film & TV Performance Games Sports Music Avg Unified-IO-2 L [42] Audio 23.0 25.4 24.2 26.7 27.7 23.7 25.0 27.1 25.2 + Caption 21.5 21.1 20.7 17.1 19.9 19.0 22.9 23.7 20.9-4.3 + Video 19.3 22.8 23.1 25.6 25.8 24.1 22.9 25.3 23.3-1.9 Unified-IO-2 XL [42] Audio 21.7 22.4 22.4 22.1 24.7 25.0 25.9 24.7 23.4 + Caption 19.9 19.8 20.8 19.2 20.2 15.9 21.7 25.5 20.7-2.7 + Video 26.5 24.4 22.5 23.5 24.7 28.0 25.7 24.2 24.7+1.3 Unified-IO-2 XXL [42] Audio 27.5 28.7 23.9 23.2 25.8 21.1 26.2 30.2 25.9 + Caption 24.0 26.7 23.0 18.9 18.7 20.7 25.9 29.4 23.7-2.2 + Video 27.1 31.7 23.9 23.7 25.5 23.7 25.7 27.3 25.9+0.0 OneLLM [21] Audio 25.7 26.1 19.3 21.9 25.8 25.9 21.5 22.4 23.0 + Caption 29.6 29.0 25.9 29.1 33.0 26.7 29.2 28.6 28.6+5.6 + Video 26.7 25.1 19.0 22.7 27.0 23.7 22.4 19.8 22.8-0.2 VideoLLaMA2 [7] Audio 23.8 23.4 21.3 22.4 24.7 19.8 27.1 27.9 23.8 + Caption 30.0 30.0 25.6 29.9 28.5 25.0 29.7 29.9 28.5+4.7 + Video 29.4 25.4 21.8 24.5 26.2 24.6 25.5 27.1 25.4+1.6 Gemini 1.5 Pro [58] Audio 40.2 42.9 35.8 33.3 33.0 31.0 33.3 24.7 34.6 + Caption 49.5 52.1 41.8 42.9 46.4 41.8 39.6 36.7 43.6+9.0 + Video 53.7 47.2 50.3 50.4 52.4 46.8 40.2 42.0 48.0+13.4
-Second and surprisingly, existing open-source audio-visual MLLMs perform even worse, achieving accuracy rates comparable to random guessing and notably below video-only MLLMs.
+Second and surprisingly, most of existing open-source audio-visual MLLMs perform even worse, achieving accuracy rates comparable to random guessing and notably below video-only MLLMs.
 This counter-intuitive finding reveals that despite having access to both modalities, these models struggle with effective audio-visual integration, suggesting that multimodal processing capability alone does not guarantee better performance without sophisticated integration mechanisms.
-Third, among proprietary MLLMs, vision-only models GPT-4o and Claude 3.5 Sonnet perform similarly to the best open-source video MLLMs.
+Third, among proprietary MLLMs, vision-only models GPT-4o and Claude 3.5 Sonnet demonstrate performance comparable to the leading open-source video MLLMs.
 Gemini 1.5 Pro, capable of processing both audio and visual information, achieves the highest accuracy of .
@@
 These findings point to crucial directions for future research and development in MLLMs.
-Breakdown Results.
-We conduct a fine-grained analysis of model performance across different audio types and task categories, as shown in Figure 4, revealing important insights into the limitations of existing multimodal models.
+Breakdown Results. We conduct a fine-grained analysis of model performance across different audio types and task categories, as shown in Figure 4 and 7, highlighting the limitations of MLLMs.
 First, models consistently underperform on audio-related tasks (e.g., audio recognition, audio counting) compared to other task types, demonstrating significant challenges in audio understanding.
@@
 While Gemini 1.5 Pro performs best overall, it shows notably lower accuracy on event-related questions compared to speech or music tasks, possibly due to the complex nature of environmental sounds.
-Other models also exhibit inconsistent performance across audio types, underscoring a general limitation for existing models for audio understanding.
+Other models also exhibit inconsistent performance across audio types, underscoring a general limitation for audio understanding.
+Methods Modality Tech & Science Culture & Politics Daily Life Film & TV Perfor- mance Games Sports Music Avg Unified-IO-2 L [43] Audio 23.0 25.4 24.2 26.7 27.7 23.7 25.0 27.1 25.2 + Caption 21.5 21.1 20.7 17.1 19.9 19.0 22.9 23.7 20.9-4.3 + Video 19.3 22.8 23.1 25.6 25.8 24.1 22.9 25.3 23.3-1.9 Unified-IO-2 XL [43] Audio 21.7 22.4 22.4 22.1 24.7 25.0 25.9 24.7 23.4 + Caption 19.9 19.8 20.8 19.2 20.2 15.9 21.7 25.5 20.7-2.7 + Video 26.5 24.4 22.5 23.5 24.7 28.0 25.7 24.2 24.7+1.3 Unified-IO-2 XXL [43] Audio 27.5 28.7 23.9 23.2 25.8 21.1 26.2 30.2 25.9 + Caption 24.0 26.7 23.0 18.9 18.7 20.7 25.9 29.4 23.7-2.2 + Video 27.1 31.7 23.9 23.7 25.5 23.7 25.7 27.3 25.9+0.0 OneLLM [22] Audio 25.7 26.1 19.3 21.9 25.8 25.9 21.5 22.4 23.0 + Caption 29.6 29.0 25.9 29.1 33.0 26.7 29.2 28.6 28.6+5.6 + Video 26.7 25.1 19.0 22.7 27.0 23.7 22.4 19.8 22.8-0.2 VideoLLaMA2 [7] Audio 23.8 23.4 21.3 22.4 24.7 19.8 27.1 27.9 23.8 + Caption 30.0 30.0 25.6 29.9 28.5 25.0 29.7 29.9 28.5+4.7 + Video 29.4 25.4 21.8 24.5 26.2 24.6 25.5 27.1 25.4+1.6 VITA-1.5 [18] Audio 30.2 35.6 36.3 30.9 32.2 32.2 31.4 33.3 32.9 + Caption 39.2 39.8 37.2 37.5 37.5 35.2 34.9 38.4 37.5+4.6 + Video 38.2 35.9 34.3 39.8 41.2 32.6 34.7 39.9 36.9+4.0 Qwen2.5-Omni [71] Audio 40.0 38.2 36.0 33.5 31.1 30.5 32.3 33.3 34.9 + Caption 40.0 37.9 38.9 33.5 36.7 37.8 37.7 38.9 37.9+3.0 + Video 47.8 49.8 43.6 43.8 48.3 39.1 43.5 47.3 45.4+10.5 Gemini 1.5 Pro [60] Audio 40.2 42.9 35.8 33.3 33.0 31.0 33.3 24.7 34.6 + Caption 49.5 52.1 41.8 42.9 46.4 41.8 39.6 36.7 43.6+9.0 + Video 53.7 47.2 50.3 50.4 52.4 46.8 40.2 42.0 48.0+13.4
 4.3 Roadmap Towards Real-world Understanding
@@
 For Gemini 1.5 Pro, accuracy increases from (video-only) to with subtitles, and further to with original audio.
-OneLLM shows similar improvements.
-These results demonstrate that both subtitles and acoustic features (including tone, emotion, and environmental sounds) contribute valuable information for multimodal understanding, beyond what subtitles alone can capture, highlighting the importance of complete acoustic cues for multimodal understanding.
+Other models, such as OneLLM and Qwen2.5-Omni, show similar improvements.
+These results demonstrate that both subtitles and acoustic features (including tone, emotion, and environmental sounds) contribute valuable information, beyond what subtitles alone can capture, emphasizing the importance of complete acoustic cues in omni-modal real-world understanding.
 Interestingly, UnifiedIO2 demonstrates performance degradation when integrating either subtitles or audio, with subtitles causing a notable accuracy decline, suggesting difficulties in multimodal processing.
-Conversely, Video-LLaMA2 improves with both modalities but performs better with subtitles than original audio, indicating a stronger reliance on textual rather than complex acoustic information.
+Conversely, Video-LLaMA2 improves with both modalities but performs better with subtitles than original audio, indicating a stronger reliance on textual rather than acoustic information.
 We further evaluate video-only MLLMs by providing transcribed subtitles, as shown in Table 4.
@@
 These insights suggest important directions for improving MLLMs’ ability to integrate acoustic and textual information for comprehensive scene understanding.
-Methods Speech Event Music Overall Video + Subtitle + Audio Video + Subtitle + Audio Video + Subtitle + Audio Video + Subtitle + Audio Unified-IO-2 L [42] 26.8 13.9-12.9 23.1-3.1 26.9 13.5-13.4 23.2-3.7 26.3 15.0-11.3 24.5-1.8 26.6 14.8-11.8 23.3-3.3 Unified-IO-2 XL [42] 25.0 13.0-12.0 24.6-0.4 24.8 12.3-12.5 25.0+0.2 26.7 15.9-10.8 24.7-2.0 25.3 14.1-11.2 24.7-0.6 Unified-IO-2 XXL [42] 27.0 15.6-11.4 25.7-1.3 26.2 14.2-12.0 25.1-1.1 28.4 19.1-9.3 25.9-2.5 27.2 17.2-10.0 25.9-1.3 OneLLM [21] 12.5 19.6+7.1 23.2+10.7 12.4 19.3+6.9 22.3+9.9 12.4 19.0+6.6 22.5+10.1 12.6 19.6+7.0 22.8+10.2 VideoLLaMA2 [7] 17.1 25.5+8.4 25.2+8.1 16.1 24.9+8.8 24.5+8.4 17.7 27.0+9.3 25.6+7.9 17.4 26.1+8.7 25.4+8.0 Gemini 1.5 Pro [58] 34.3 39.6+5.3 49.2+14.9 33.0 38.9+5.9 48.7+15.7 35.4 39.2+3.8 49.8+14.4 34.4 39.3+4.9 48.0+13.6
+Methods Speech Event Music Overall Video + Subtitle + Audio Video + Subtitle + Audio Video + Subtitle + Audio Video + Subtitle + Audio Unified-IO-2 L [43] 26.8 13.9-12.9 23.1-3.1 26.9 13.5-13.4 23.2-3.7 26.3 15.0-11.3 24.5-1.8 26.6 14.8-11.8 23.3-3.3 Unified-IO-2 XL [43] 25.0 13.0-12.0 24.6-0.4 24.8 12.3-12.5 25.0+0.2 26.7 15.9-10.8 24.7-2.0 25.3 14.1-11.2 24.7-0.6 Unified-IO-2 XXL [43] 27.0 15.6-11.4 25.7-1.3 26.2 14.2-12.0 25.1-1.1 28.4 19.1-9.3 25.9-2.5 27.2 17.2-10.0 25.9-1.3 OneLLM [22] 12.5 19.6+7.1 23.2+10.7 12.4 19.3+6.9 22.3+9.9 12.4 19.0+6.6 22.5+10.1 12.6 19.6+7.0 22.8+10.2 VideoLLaMA2 [7] 17.1 25.5+8.4 25.2+8.1 16.1 24.9+8.8 24.5+8.4 17.7 27.0+9.3 25.6+7.9 17.4 26.1+8.7 25.4+8.0 VITA-1.5 [18] 37.6 39.1+1.5 36.2-1.4 36.4 38.2+1.8 35.3-1.1 38.7 40.0+1.3 37.1-1.6 37.7 39.3+1.6 36.5-1.2 Qwen2.5-Omni [71] 38.7 38.7+0.0 44.8+6.1 37.6 37.7+0.1 43.8+6.2 40.7 40.3-0.4 46.1+5.4 39.2 39.2+0.0 45.2+6.0 Gemini 1.5 Pro [60] 34.3 39.6+5.3 49.2+14.9 33.0 38.9+5.9 48.7+15.7 35.4 39.2+3.8 49.8+14.4 34.4 39.3+4.9 48.0+13.6
 Methods Speech Event Music Overall Video + Subtitle Video + Subtitle Video + Subtitle Video + Subtitle Video-LLaVA [33] 20.3 15.4-4.9 19.8 14.4 -5.4 19.5 16.4-3.1 20.3 16.0-4.3 LLaMA3.2 [21] 27.1 29.3+2.2 27.6 29.6+2.0 25.9 28.1+2.2 27.1 28.8+1.7 Qwen2-VL [64] 31.8 41.1+9.3 30.9 39.4+8.5 34.2 41.8+7.6 32.4 41.2+8.8 mPLUG-Owl3 [75] 33.0 39.2+6.2 32.3 38.3+6.0 34.6 39.2+4.6 32.9 38.7+5.8 LLaVA-OneVision [27] 37.7 44.0+6.3 36.3 42.7+6.4 39.7 45.7+6.0 37.7 43.9+6.2 InternVL2.5 [5] 39.0 48.3+9.3 38.6 47.9+9.3 39.2 47.1+7.9 39.1 47.8+8.7 LLaVA-Video [82] 40.5 45.9+5.4 38.9 44.6+5.7 42.3 47.7+5.4 40.2 45.6+5.4 GPT 4o [25] 42.8 51.1+8.3 40.9 50.2+9.3 43.6 49.9+6.3 42.6 50.1+7.5
-Video Frames.
-We investigate the effect of temporal sampling density by varying input frame numbers for video-only MLLMs.
-As shown in Figure 5, most models demonstrate significant performance improvements with increased frame density, with LLaMA-3.2 being a notable exception.
-These gains likely result from better capture of fine-grained temporal dynamics and subtle visual changes, emphasizing the importance of dense temporal sampling.
+Failure Analysis and Future Improvement.
+We perform error analysis on 130 samples of Gemini 1.5 Pro (5 random samples per task) through manual review, identifying three main error types: Audio Understanding Errors (misinterpreting audio information), Visual Understanding Errors (missing visual details), and Reasoning Errors (faulty logical steps).
+As shown in Figure 7, most errors stem from audio understanding deficiencies and reasoning failures.
+The reason for poor accuracy and limitation of existing models can be summarized as follows: (i) Inadequate Audio Understanding.
+Existing models fail to understand audio information correctly and show significantly weaker audio processing than visual understanding.
+(ii) Limited Cross-Modal Integration.
+Models often process modalities independently rather than performing true multimodal integration and suffer from insufficient omni-modal information integration.
+(iii) Insufficient Complex Reasoning Ability.
+Despite correct perception, MLLMs still conduct error reasoning, leading to incorrect conclusions.
+We also raise several key strategies to enhance models’ understanding of omni-modality information: (i) Coupled Multimodal Training Data.
+Using naturally coupled, interleaved multimodal data, for example, audio, visual, language content, would enhance models’ capability to leverage cross-modal dependencies.
+(ii) Architectural Improvements.
+Enhanced attention mechanisms facilitating deep multimodal integration could emphasize early fusion between modalities, rather than processing them as separate streams for late fusion.
+(iii) Advanced Modal Alignment Techniques.
+Progressive alignment strategies that gradually enhance the model’s ability to align information across modalities could lead to more effective utilization of multimodal inputs.
+(iiii) Reasoning strengthening.
+Incorporating diverse reasoning-focused data can strengthen logical inference capabilities, enabling more coherent and accurate conclusions.
 5 Conclusion
-In this paper, we introduce WorldSense, the first benchmark designed to evaluate MLLMs’ omnimodal understanding in real-world scenarios.
+In this paper, we propose WorldSense, the first benchmark designed to evaluate MLLMs’ omnimodal understanding in real-world scenarios.
 Distinguished by its emphasis on joint omnimodal comprehension across diverse real-world contexts, WorldSense encompasses rich video categories and carefully curated question-answer pairs that necessitate the integration of visual and acoustic information.
@@
 We hope that WorldSense can serve as a foundational benchmark for advancing human-like omnimodal understanding capabilities.
-Appendix A Broader Impacts
+Appendix A Quality Control
+Experienced annotators.
+Our annotation team consists of 80 professional annotators with extensive QA annotation experience.
+These annotators are proficient in English, and have participated in sereval QA data annotation projects.
+Sufficient annotation training. We conducted a one-week training program with 200 videos (excluded from final benchmark) until annotators achieved high proficiency (only requiring modifications).
+Annotation instruction.
+Each annotator received a comprehensive instruction with task explanations, question formulation guidelines, QA creation instructions, annotated examples, and cross-modal inference requirements.
+Our review process identified and revised similar QA pairs. Through professional annotators, thorough training, detailed guidelines, and rigorous quality control, we ensure high-quality annotations.
+Appendix B Implement Details
+For open-source MLLMs, we strictly follow their official implementations and recommended preprocessing pipelines to ensure fair comparison.
+For GPT 4o and Claude 3.5 Sonnet, we sample 16 frames uniformly from each video, while for Gemini 1.5 Pro, we utilize the official API for raw video file uploads.
+We conduct all the experiments on a NVIDIA A100 GPU.
+Appendix C Evaluation Prompt
+Following previous works [17, 31], we adopt the format of “whole video frames + whole subtitles/audios (optional) + question with prompt” as prompt.
+We show the evaluation prompt across three input configurations: video-only input, video with subtitles, and video with audio content as following.
+Prompt Avg Prompt 1 24.0 Prompt 2 24.8 Prompt Ours 25.4
+Appendix D Prompt Sensitivity
+We have conducted on experiments on VideoLLaMA2 to explore the impact of different prompts, and results are shown in Table 5.
+Prompt 1 stands for "Based on your observations, select the best option that accurately addresses the question.", Prompt 2 is "Carefully watch this video and pay attention to every detail.", and Prompt Ours represents the prompt in Section C.
+As shown in the experiment, the prompt template we ultimately used consistently yielded superior performance.
+Appendix E Failure Case
+Figure 8 showcases two failure cases of Gemini 1.5 Pro, VideoLLaMA2, and LLaVA-Video.
+The left case involves a vision understanding and reasoning error, where Gemini 1.5 Pro incorrectly identified the clock time as 10:30 instead of the actual 10:00 displayed, leading to the incorrect answer.
+This reflects deficiencies in basic visual perception and properly correlating visual information with the question.
+The right case demonstrates an audio understanding error, where models misjudged the rhythm pattern change in guzheng music, interpreting it as changing from soothing to intense (option D) rather than the correct passionate to soothing (option B).
+This case indicates that tasks involving interpretation of musical emotion and rhythm patterns remain challenging for existing MLLMs.
+These cases highlight the current limitations of MLLMs in processing subtle visual details and interpreting subjective audio qualities.
+Appendix F Limitation
+While our WorldSense represents a significant advancement in evaluating multimodal understanding capabilities of MLLMs, the multiple-choice format inevitably constrains the assessment of models’ generative capabilities.
+Real-world understanding often requires open-ended responses, explanations, and adaptability beyond selecting from predefined options.
+Our WorldSense may not adequately evaluate how models perform on tasks requiring nuanced reasoning or creative problem-solving.
+We will add open-ended questions and expand the evaluation paradigm to better assess real-world multimodal understanding.
+Appendix G Broader Impacts & Ethics Statement
 Our work on WorldSense has several potential positive impacts on society and AI development, while also presenting certain risks that warrant careful consideration.
@@
 Advanced multimodal understanding capabilities could potentially be misused for surveillance or monitoring purposes.
-We believe that open discussion of these impacts is crucial for responsible development of MLLMs.
-Appendix B Ethics Statement
-Our research on WorldSense adheres to strict ethical principles and guidelines. We acknowledge several important ethical considerations:
-Data Collection and Privacy.
+We believe that open discussion of these impacts is crucial for the responsible development of multi-modal large language models.
+Our research on WorldSense adheres to strict ethical principles and guidelines.
+We acknowledge several important ethical considerations: (1) Data Collection and Privacy.
 All video content in WorldSense has been collected from publicly available sources with appropriate licensing agreements.
 We have con ducted thorough reviews and implemented comprehensive data processing procedures to ensure privacy protection, including the removal of any personally identifiable information.
-Potential Biases.
+(2) Potential Biases.
 While acknowledging that inherent biases may exist in any dataset, we have undertaken systematic efforts to ensure diverse representation across our video content and question-answer pairs, encompassing various domains, cultures, and contexts.
 Nevertheless, we recognize that completely eliminating bias remains a significant challenge, and users should carefully consider these potential limitations when utilizing our dataset.
-Intended Use.
+(3) Intended Use.
 WorldSense is specifically designed to advance research in omnimodal real-world understanding.
@@
 Users are expected to adhere to ethical guidelines and responsible practices.
-Appendix C License
+Appendix H License
 The WorldSense dataset is released under the CC BY-NC-SA 4.0 License. Authors bear all responsibility in case of violation of rights and confirmation of the data license.
+Appendix I Datasheets
+I.1 Motivation
+For what purpose was the dataset created?
+To evaluate MLLMs’ capabilities in real-world omnimodal understanding.
+Who created the dataset (e.g., which team, research group) and on behalf of which entity (e.g., company, institution, organization)?
+The authors of this paper.
+Who funded the creation of the dataset?
+Xiaohongshu Inc.
+Any other comments?
+No
+I.2 Composition
+What do the instances that comprise the dataset represent (e.g., documents, photos, people, countries)?
+Videos along with captions and question/answer pairs.
+How many instances are there in total (of each type, if appropriate)?
+WorldSense contains 3,172 question-answer pairs and contains 1,662 videos in total.
+Does the dataset contain all possible instances or is it a sample (not necessarily random) of instances from a larger set?
+Videos of WorldSense are sampled from FineVideo and Music AVQA. All QA pairs are re-annotated manually.
+What data does each instance consist of?
+Each instance contains one video with its corresponding audio, a question about the video content and the corresponding answer, the category of the video, the fine-grained video understanding capability examined by the question, and the class of audio content.
+Each instance also contain the auto-generated subtitles sourced from YouTube.
+Is there a label or target associated with each instance?
+Yes. We provide the ground-truth answer for each question.
+Is any information missing from individual instances?
+N/A.
+Are relationships between individual instances made explicit (e.g., users’ movie ratings, social network links)?
+N/A.
+Are there recommended data splits (e.g., training, development/validation, testing)?
+No, WorldSense is designed for evaluation only.
+Are there any errors, sources of noise, or redundancies in the dataset?
+No.
+Is the dataset self-contained, or does it link to or otherwise rely on external resources (e.g., websites, tweets, other datasets)?
+WorldSense is self-contained.
+Does the dataset contain data that might be considered confidential (e.g., data that is protected by legal privilege or by doctor – patient confidentiality, data that includes the content of individuals’ non-public communications)?
+N/A.
+Does the dataset contain data that, if viewed directly, might be offensive, insulting, threatening, or might otherwise cause anxiety?
+N/A.
+I.3 Collection Process
+How was the data associated with each instance acquired?
+See main paper for details.
+What mechanisms or procedures were used to collect the data (e.g., hardware apparatuses or sensors, manual human curation, software programs, software APIs)?
+Humans are required to propose a question and corresponding answer based on the video. MLLMs, such as Qwen2-VL, Video-LLaMA2 and OneLLM are utilized to perform quality control.
+If the dataset is a sample from a larger set, what was the sampling strategy (e.g., deterministic, probabilistic with specific sampling probabilities)?
+Yes, we sample the videos from FineVideo and Music-AVQA. See main paper for details.
+Who was involved in the data collection process (e.g., students, crowdworkers, contractors) and how were they compensated (e.g., how much were crowdworkers paid)?
+The authors and contractors are involved in the data collection process and are paid a fair wage.
+Over what timeframe was the data collected?
+The dataset is collected in 2024.
+Were any ethical review processes conducted (e.g., by an institutional review board)?
+All videos in our benchmark are human-selected based on appropriate value propositions and undergo a second manual quality check to ensure there are no ethical violations.
+Did you collect the data from the individuals in question directly, or obtain it via third parties or other sources (e.g., websites)?
+We obtained video data from FineVideo and Music-AVQA.
+Were the individuals in question notified about the data collection?
+We didn’t collect the data from the individuals. The data was collected from public web sources instead.
+Did the individuals in question consent to the collection and use of their data?
+N/A.
+If consent was obtained, were the consenting individuals provided with a mechanism to revoke their consent in the future or for certain uses?
+N/A.
+Has an analysis of the potential impact of the dataset and its use on data subjects (e.g., a data protection impact analysis) been conducted?
+N/A.
+Any other comments?
+No.
+I.4 Preprocessing/cleaning/labeling
+Was any preprocessing/cleaning/labeling of the data done (e.g., discretization or bucketing, tokenization, part-of-speech tagging, SIFT feature extraction, removal of instances, processing of missing values)?
+We firstly select videos based on pre-designed categories, and then clip the video based on visual-audio correlation and dynamic scores.
+Was the “raw” data saved in addition to the preprocessed/cleaned/labeled data (e.g., to support unanticipated future uses)?
+N/A.
+Is the software that was used to preprocess/clean/label the data available?
+We use the open-source models.
+Any other comments?
+No.
+I.5 Uses
+Has the dataset been used for any tasks already?
+Yes. We have used the dataset to evaluate video question answering in real-world.
+Is there a repository that links to any or all papers or systems that use the dataset?
+No.
+What (other) tasks could the dataset be used for?
+It also can be used to evaluate the video understanding capability of VLMs.
+Is there anything about the composition of the dataset or the way it was collected and preprocessed/cleaned/labeled that might impact future uses?
+No.
+Are there tasks for which the dataset should not be used?
+N/A.
+Any other comments?
+No.
+I.6 Distribution
+Will the dataset be distributed to third parties outside of the entity (e.g., company, institution, organization) on behalf of which the dataset was created?
+Yes, the dataset will be made publicly available.
+How will the dataset will be distributed (e.g., tarball on website, API, GitHub)?
+We host it on the webpage, GitHub, and Huggingface.
+When will the dataset be distributed?
+It’s availale and open to the public now.
+Will the dataset be distributed under a copyright or other intellectual property (IP) license, and/or under applicable terms of use (ToU)?
+We release our benchmark under CC BY-NC 4.0 license.
+Have any third parties imposed IP-based or other restrictions on the data associated with the instances?
+No.
+Do any export controls or other regulatory restrictions apply to the dataset or to individual instances?
+No.
+Any other comments?
+No.
+I.7 Maintenance
+Who will be supporting/hosting/maintaining the dataset?
+The authors will be supporting/hosting/maintaining the dataset.
+How can the owner/curator/manager of the dataset be contacted (e.g., email address)?
+No.
+Is there an erratum?
+Currently, we do not have an erratum. We will update if we find errors.
+Will the dataset be updated (e.g., to correct labeling errors, add new instances, delete instances)?
+Yes. We will make announcements on GitHub if there is any update.
+If the dataset relates to people, are there applicable limits on the retention of the data associated with the instances (e.g., were the individuals in question told that their data would be retained for a fixed period of time and then deleted)?
+N/A.
+Will older versions of the dataset continue to be supported/hosted/maintained?
+Yes.
+If others want to extend/augment/build on/contribute to the dataset, is there a mechanism for them to do so?
+Yes. Contributors can post issues or submit pull requests on GitHub. We will review and verify contributions, and update the dataset if the contribution is useful.
+Any other comments?
+No.